Hello out there,I need some help with some square root problems,the first one is,Estimate the value of the square root:(it has a square root symbol above it)0.85.The next one is,If x=(square root symbol)0.04,what is the value of x?The next one is,Simplify:-(square root symbol)0.01 divided by 0.25.And the last one is,Simplify :(square root symbol)36 divided by 225.Thanks to whoever can help me.
Answer by praseenakos@yahoo.com(507)


square root of 0.85=0.21 (by division method0
if x= square root of 0.04 then x= 0.2
-(square root of 0.01)/(square root of 0.25)
= -0.1/0.5 = -1/5